### v1

@@ 5 T_POS pVal;
In our experiments, we test mutants in parallel, by leveraging the multiple nodes of an HPC infrastructure HPC . This decision depends on the observation that leveraging parallelism (e.g., by relying on Cloud solutions) is a cost-effective choice in the case of complex CPS with many live mutants. Parallel testing of mutants is enabled by the fact that, in contrast to from MU2 MU2 , MOTIF does not test live mutants with the same inputs. Therefore, MOTIF can simply be deployed on different nodes and executed. However, we assess (RQ4) how MOTIF’s effectiveness can be improved by reusing the mutant-killing inputs obtained in the different, parallel executions.
@@ 5.3.2 Results
The compilers’ comparison results show that Clang outperforms GCC by 1.37 percentage points (pp) in LIBU, 1.13 pp in ASN1lib, 2.91 pp in MLFS, and 0.65 pp in ESAIL. For each subject of these four subjects, the results obtained with the two compilers are significantly different at every minute. The better performance of Clang is likely due to the fact that the number of inputs generated by AFL++ with Clang is 4 to 20 times higher than the one observed with GCC. Indeed, search algorithms (i.e., AFL++) are more likely to explore a larger portion of the input space when they generate more inputs. The larger number of inputs generated with Clang can be explained by Clang generating executable programs (in our case, fuzz drivers) that are quicker to execute than GCC ones and therefore AFL++ can try more inputs with Clang than with GCC, for the same test budget. In contrast to the above, for S5, Clang performed worse than GCC by 0.30 pp but the difference is not statistically significant, enabling us to conclude that overall Clang is the best choice for MOTIF.
@@
The primary reason why MOTIF with sanitizers cannot outperform the baseline is the overhead of the instrumented code introduced by sanitizers to detect invalid behaviors. Such overhead reduces the number of inputs generated by AFL++, consequently decreasing the number of mutants killed, even though sanitizers may kill some mutants not killed without them. In our investigation, the worst performance is obtained with MSAN, which led to 96.77% fewer inputs than the baseline, on average, and showed significantly lower performance. The worst MSAN result is obtained with S5, likely because of poor support for C++ libraries (e.g., MSAN led to crashes whenever it encounters a standard library function such as std::map and std::set). ASAN, instead, led to 61.5% fewer inputs, on average, but enabled the detection of a large number of additional mutants thus compensating the loss due to execution cost; consequently, ASAN achieved results that are comparable to those obtained without sanitizer. ASAN effectiveness in killing mutants mainly depends on its capability to detect out-of-bound memory accesses caused by mutants; such mutants are harder to detect otherwise because the sufficiency condition (see Section 2.3) is hard to meet without a sanitizer (the effect of the out-of-bound access shall affect the values assigned to an output variable). Further, ASAN prevents the generation of test inputs that kill mutants but are not valid; those are inputs that violate function preconditions and cause out-of-bound accesses in the original function under test, and are thus discarded by MOTIF. Different from ASAN, UBSAN prevents the testing of several mutants because it detects invalid behaviors in the original function (on average, 653.8 fuzzing drivers failed when executing the original function), which need to be fixed before proceeding with mutation testing. A representative example is provided by MLFS, where UBSAN detects several arithmetic overflows caused by shift operators applied to variables that may have negative values. Although UBSAN may have detected potential bugs in the SUT, by terminating the execution of the original function, UBSAN prevents mutation testing. In summary, ASAN is the only sanitizer that is beneficial to MOTIF for two of the subjects (i.e., prevents false positives without largely affecting effectiveness).
@@
Although our results did not help identifying a sanitizer and coverage metric configuration that provides effectiveness improvements across all subjects, for the sake of simplifying the remaining experiments, we identify a configuration for MOTIF that is likely to provide improvements in most subjects. It consists of combining Clang, ASAN, and LAF and we call such configuration Best, . We selected ASAN because it performs similarly to the baseline in terms of percentage of killed mutants but also prevents the generation of inputs that are invalid and cause memory access errors. We selected LAF because it leads to the largest number of killed mutants, across all subjects (3283 versus 3186 for NGRAM and 3190 for CTX), and further leads to the largest effectiveness improvement (+2.27 pp), and also improves the subject with the most effective test suite (i.e., MLFS, which achieves MC/DC and 81.80% MS).
@@
To ensure that the selected configuration options do not interfere, we performed ten additional executions of MOTIF configured with the Best settings, with each subject, and compare the observed results with MOTIF relying on the default AFL++ options (i.e., relying on GCC without additional sanitizers or coverage metrics). Figure 3 shows the percentage of mutants killed by MOTIF after each second, when using GCC and Best, for each subject. Each line captures the average percentage observed in ten runs, with the shaded area capturing the upper and lower bounds across those runs. The vertical dashed line shows the time budget for the experiment. With Best, the percentages of killed mutants are 49.84% for LIBU, 88.74% for ASN1lib, 39.85% for MLFS, 39.50% for ESAIL, and 82.53% for S5. These performance results are significantly better than those of our baseline (i.e., MOTIF with GCC) for four subjects: it increases the percentage of killed mutants by 1.53 pp in LIBU, 3.30 pp in ASN1lib, 4.34 pp in MLFS, and 1.12 pp in ESAIL. For S5, the performance is similar to that of the baseline. Further, we manually inspected the mutants killed only by the Best configuration and observed that 70% of the generated test cases enable exercising boundary cases missed by the original test suite, which include access to first/last items of arrays and use of boundary values in conditional statement. We can thus conclude that Best helps improve the quality of test suites because in safety-critical systems it is desirable to exercise such situations. We can thus conclude that combining Clang, ASAN, and LAF leads to the best results in MOTIF.
@@ 5.5.2 Results
Concluding, although the selected seed inputs help kill mutants, the contribution of the fuzzing process is significant with, at the very minimum (S5), 52.27% (i.e., $100\%-\frac{39.39\%/}{82.53\%}$) of the killed mutants being killed by fuzzing.
@@ 5.6.2 Results
As for the time taken to kill mutants, we report that the mutants in ASN1lib and MLFS were killed significantly faster with MOTIF-Hybrid. The median per mutant for ASN1lib and for MLFS is 142.00 vs. 208.66 seconds and 214.60 vs. 253.08 seconds, respectively, when compared with MOTIF. In LIBU and ESAIL, the difference is not significant, however, with the median observed for MOTIF-Hybrid with LIBU being lower (57.13 seconds with MOTIF-Hybrid vs. 66.13 seconds with MOTIF).
@@ 5.7 RQ5 - Reusing inputs
Table 6: Comparison MOTIF with Best set of fuzzing options and results after reusing inputs for five selected subjects
Subject Target mutants 3 c # of killed mutants (Avg. of 10 runs) 3 c # of inputs for functions (10 runs) 3 c Reuse time (50 runs, seconds) 3 c Kill time (50 runs, seconds)
@@
n.nn pp%: outperformed the baseline with $p\leq 0.01$,
5.7.1 Design
As for cost, we measure, for each mutant, the time taken to execute fuzzing drivers with all inputs and the time taken until the mutant is killed. Since the order of inputs may affect execution time, we repeat our experiment 50 times, after shuffling each input set.
5.7.2 Results
Table 5.7 presents our results. Column target mutants reports the number of mutants that belong to the functions that have at least one killed mutant. Column FUZZED reports the MS obtained by MOTIF on average over 10 runs. Column REUSED reports the MS obtained after reusing inputs to kill additional mutants. Column DIFF shows the differences between FUZZED and REUSED, and reports on the significance of the difference (we highlight significant improvements). The next three sets of columns report statistics (mean, min, and max) on the number of inputs available for reuse across target functions, the time required to test each mutant with the reused inputs, and the time taken to kill a mutant.
The experiment results indicate that MS can be improved by up to 3.79 pp when reusing inputs, with significant improvements in most of the subjects, which indicates that input reuse is beneficial for MOTIF. The time taken for reusing inputs varies, taking up to 15 minutes (900 seconds) depending on several factors, including the number of inputs, software size, and the type of mutants. For instance, ESAIL has approximately 70% of its mutants within a single function, which leads to a large number of additional executions when reusing inputs. In the case of S5, the larger size causes a longer execution time for fuzzing drivers compared to other subjects, but the maximum execution time is lower due to the number of inputs collected. The execution time is also affected by the type of mutants. Specifically, some mutants took longer than others because they modified stopping conditions in program loops.
@@ 5.7.2 Results
5.8 Threats to validity
@@
Though our results may depend on the specific fuzzer used in our experiments, AFL++ is one of the best performing grey-box fuzzers according to recent benchmarks (see Section 5.2). Further, we assessed MOTIF with a state-of-the art hybrid fuzzing solution. Although alternative hybrid fuzzers were not applicable to some of our subjects (see Section 5.2), they might have led to different and potentially better results; we leave their investigation to future work.
@@
To address generalizability issues, we selected diverse software subjects that are installed and running on space CPS, including satellites currently in orbit: a mathematical library, a utility library, a data serialization component, onboard control software, and ground software. Since they implement a diverse set of features (mathematical operations, serialization, string and time utilities), they strengthen the generalizability of our results. Further, these types of software components are typical in many CPS systems including avionics, robotics, and automotive, thus suggesting the proposed approach may be useful in many sectors other than space.

### v2

@@ 5 T_POS pVal;
In our experiments, we test mutants in parallel, by leveraging the multiple nodes of an HPC infrastructure HPC . This decision depends on the observation that leveraging parallelism (e.g., by relying on Cloud solutions) is a cost-effective choice in the case of complex CPS with many live mutants. Parallel testing of mutants is enabled by the fact that, in contrast to MU2 MU2 , MOTIF does not test live mutants with the same inputs. Therefore, MOTIF can simply be deployed on different nodes and executed. However, we assess (RQ4) how MOTIF’s effectiveness can be improved by reusing the mutant-killing inputs obtained in the different, parallel executions.
@@
In the following, we discuss the test effectiveness of each tool in terms of percentage of live mutants killed by the test cases generated by the tool. Further, we compare the tools’ effectiveness in terms of difference in percentage (pp) of killed mutants:
[…]
For example, with $MS_{A}=89.1$ and $MS_{B}=87.2$, we report that $A$ outperforms $B$ by 1.9 pp.
@@ 5.3.1 Design
Last, to discuss complementarity across configurations, we compare the number of mutants uniquely killed or missed by each configuration.
@@ 5.3.2 Results
The compilers’ comparison results show that Clang outperforms GCC by 1.37 pp in LIBU, 1.13 pp in ASN1lib, 2.91 pp in MLFS, and 0.65 pp in ESAIL. For each of these four subjects, the results obtained with the two compilers are significantly different at every minute. The better performance of Clang is likely due to the fact that the number of inputs generated by AFL++ with Clang is 4 to 20 times higher than the one observed with GCC. Indeed, search algorithms (i.e., AFL++) are more likely to explore a larger portion of the input space when they generate more inputs. The larger number of inputs generated with Clang can be explained by Clang generating executable programs (in our case, fuzz drivers) that are quicker to execute than GCC ones and therefore AFL++ can try more inputs with Clang than with GCC, for the same test budget. In contrast to the above, for S5, Clang performed worse than GCC by 0.30 pp but the difference is not statistically significant, enabling us to conclude that overall Clang is the best choice for MOTIF.
@@
The primary reason why MOTIF with sanitizers cannot outperform the baseline is the overhead of the instrumented code introduced by sanitizers to detect invalid behaviors. Such overhead reduces the number of inputs generated by AFL++, consequently decreasing the number of mutants killed, even though sanitizers may kill some mutants not killed without them. In our investigation, the worst performance is obtained with MSAN, which led to 96.77% fewer inputs than the baseline, on average, and showed significantly lower performance. The worst MSAN result is obtained with S5, likely because of poor support for C++ libraries (e.g., MSAN led to crashes whenever it encounters a standard library function such as std::map and std::set). ASAN, instead, led to 61.5% fewer inputs, on average, but enabled the detection of a large number of additional mutants thus compensating the loss due to execution cost; consequently, ASAN achieved results that are comparable to those obtained without sanitizer. ASAN effectiveness in killing mutants mainly depends on its capability to detect out-of-bound memory accesses caused by mutants; such mutants are harder to detect otherwise because the sufficiency condition (see Section 2.3) is hard to meet without a sanitizer (the effect of the out-of-bound access shall affect the values assigned to an output variable). Further, ASAN prevents the generation of test inputs that kill mutants but are not valid; those are inputs that violate function preconditions and cause out-of-bound accesses in the original function under test, and are thus discarded by MOTIF. Different from ASAN, UBSAN prevents the testing of several mutants because it detects invalid behaviors in the original function (on average, 653.8 fuzzing drivers failed when executing the original function), which need to be fixed before proceeding with mutation testing. A representative example is provided by MLFS, where UBSAN detects several arithmetic overflows caused by shift operators applied to variables that may have negative values; MLFS developers reported that UBSAN findings are overall useful for software development (e.g., they considered them to avoid mistakes when refactoring the library for a newer version, called LibmCS libMCS ) but they did not detect any fault (i.e., the overflows are not possible because of implicit function preconditions). Although UBSAN may have detected potential bugs in the SUT, by terminating the execution of the original function, UBSAN prevents mutation testing. For critical software, we therefore suggest using slightly modified versions of MOTIF fuzz drivers (e.g., invoking only the function under test), to identify possible bugs with UBSAN; however, assessing the effectiveness of this approach is out of the scope of this paper. In summary, ASAN is the only sanitizer that is beneficial to MOTIF for two of the subjects (i.e., prevents false positives without largely affecting effectiveness).
@@
Although our results did not help identifying a sanitizer and coverage metric configuration that provides effectiveness improvements across all subjects, for the sake of simplifying the remaining experiments, we identify a configuration for MOTIF that is likely to provide improvements in most subjects. It consists of combining Clang, ASAN, and LAF and we call such configuration Best. We selected ASAN because it performs similarly to the baseline in terms of percentage of killed mutants but also prevents the generation of inputs that are invalid and cause memory access errors. We selected LAF because it leads to the largest number of killed mutants, across all subjects (3283 versus 3186 for NGRAM and 3190 for CTX), and further leads to the largest effectiveness improvement (+2.27 pp), and also improves the subject with the most effective test suite (i.e., MLFS, which achieves MC/DC and 81.80% MS).
@@
To ensure that the selected configuration options do not interfere, we performed ten additional executions of MOTIF configured with the Best settings, with each subject, and compare the observed results with MOTIF relying on the default AFL++ options (i.e., relying on GCC without additional sanitizers or coverage metrics). Figure 3 shows the percentage of mutants killed by MOTIF after each second, when using GCC and Best, for each subject. Each line captures the average percentage observed in ten runs, with the shaded area capturing the upper and lower bounds across those runs. The vertical dashed line shows the time budget for the experiment. With Best, the percentages of killed mutants are 49.84% for LIBU, 88.74% for ASN1lib, 39.85% for MLFS, 39.50% for ESAIL, and 82.53% for S5. These performance results are significantly better than those of our baseline (i.e., MOTIF with GCC) for four subjects: it increases the percentage of killed mutants by 1.53 pp in LIBU, 3.30 pp in ASN1lib, 4.34 pp in MLFS, and 1.12 pp in ESAIL. For S5, the performance is similar to that of the baseline. Further, we manually inspected the mutants killed only by the Best configuration and observed that 70% of the generated test cases exercise boundary cases missed by the original test suite, which include access to first/last items of arrays and use of boundary values in conditional statement. We can thus conclude that Best helps improve the quality of test suites because, in safety-critical systems, it is desirable to exercise such situations. We can thus conclude that combining Clang, ASAN, and LAF leads to the best results in MOTIF.
@@
Table 7: Uniquely killed mutants per MOTIF configuration.
LIBU ASN1lib MLFS ESAIL S5
# of mutants 437 1347 3891 581 99
GCC 1 (0.23%) 0 (0.00%) 7 (0.18%) 0 (0.00%) 0 (0.00%)
CLANG 1 (0.23%) 0 (0.00%) 6 (0.15%) 0 (0.00%) 0 (0.00%)
ASAN 1 (0.23%) 5 (0.37%) 8 (0.21%) 4 (0.69%) 1 (1.01%)
UBSAN 4 (0.92%) 0 (0.00%) 91 (2.34%) 0 (0.00%) 0 (0.00%)
LSAN 0 (0.00%) 0 (0.00%) 2 (0.05%) 0 (0.00%) 0 (0.00%)
MSAN 2 (0.46%) 0 (0.00%) 0 (0.00%) 1 (0.17%) 0 (0.00%)
LAF 2 (0.46%) 16 (1.19%) 45 (1.16%) 0 (0.00%) 0 (0.00%)
NGRAM 1 (0.23%) 0 (0.00%) 4 (0.10%) 0 (0.00%) 0 (0.00%)
CTX 1 (0.23%) 0 (0.00%) 3 (0.08%) 0 (0.00%) 0 (0.00%)
Table 8: Uniquely missed mutants per MOTIF configuration.
LIBU ASN1lib MLFS ESAIL S5
# of mutants 437 1347 3891 581 99
GCC 2 (0.46%) 1 (0.07%) 19 (0.49%) 1 (0.17%) 0 (0.00%)
CLANG 4 (0.92%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
ASAN 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
UBSAN 0 (0.00%) 5 (0.37%) 500 (12.85%) 0 (0.00%) 0 (0.00%)
LSAN 8 (1.83%) 0 (0.00%) 1 (0.03%) 1 (0.17%) 0 (0.00%)
MSAN 9 (2.06%) 9 (0.67%) 9 (0.23%) 4 (0.69%) 75 (75.76%)
LAF 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
NGRAM 0 (0.00%) 1 (0.07%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
CTX 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Complementarity. Table 7 reports the number of mutants uniquely killed by each MOTIF configuration; they are obtained by subtracting, from the set of mutants killed by one configuration in at least one run, the set of mutants killed by the other configurations in at least one run. Table 8 presents the number of mutants missed by each configuration; they are obtained by subtracting the set of mutants killed by one configuration from the set of mutants killed by all the other configurations. Except for MOTIF with ASAN, MOTIF with UBSAN, and MOTIF with LAF, all the other MOTIF configurations kill a limited number of unique mutants (Table 7). Specifically, MOTIF with UBSAN and MOTIF with LAF are the only configurations killing more than 50 unique mutants; MOTIF with ASAN is the only configuration killing at least one unique mutant in each subject. Looking at the uniquely missed mutants (Table 8), MOTIF with GCC, MOTIF with UBSAN, and MOTIF with MSAN are the configurations leading to the highest number of missed mutants; indeed, each of them leads to more than 20 uniquely missed mutants, the worst case (505 uniquely missed mutants) being MOTIF with UBSAN.
To summarize, MOTIF with UBSAN is complementary to other configurations (95 unique mutants killed) but can’t be used as the only sanitizer (500 missed mutants); engineers may use it in addition to other configurations, if test budget is available. Also, Table 7 and 8 confirm the effectiveness of the Best configuration. Indeed, the other best performing configuration options (ASAN and LAF) are part of the Best configuration; further, none of the worst configuration options (i.e., GCC, UBSAN, MSAN) are part of Best.
@@ 5.5.2 Results
Concluding, although the selected seed inputs help kill mutants, the contribution of the fuzzing process is significant with, at the very minimum (S5), 52.27% (i.e., $100\%-\frac{39.39\%}{82.53\%}$) of the killed mutants being killed by fuzzing.
@@ 5.6.2 Results
As for the time taken to kill mutants, we report that the mutants in ASN1lib and MLFS were killed significantly faster with MOTIF-Hybrid. When compared with MOTIF, the median kill time per mutant of MOTIF-Hybrid is reduced by 31.95% for ASN1lib (from 208.66 to 142.00 seconds) and by 15.20% for MLFS (from 253.08 to 214.60 seconds). In LIBU and ESAIL, the difference is not significant, however, with the median observed for MOTIF-Hybrid with LIBU being lower by 13.65% (57.13 seconds with MOTIF-Hybrid vs. 66.13 seconds with MOTIF).
@@ 5.7 RQ5 - Reusing inputs
5.7.1 Design
As for cost, we measure, for each mutant, the time taken to execute fuzzing drivers with all inputs and the time taken until the mutant is killed. Since the order of inputs may affect execution time, we repeat our experiment 50 times, after shuffling each input set, for each of the 10 MOTIF runs for the Best configuration considered in RQ1. Our setting leads to 500 different results for execution time, in total. By construction, all the 50 repetitions lead to the same mutation score (the final mutation score is not based on the input processing order); consequently, we have 10 distinct mutation score results, not 500.
5.7.2 Results
Table 5.7.2 presents our results. Column target mutants reports the number of mutants that belong to the functions that have at least one killed mutant. Column FUZZED reports the MS obtained by MOTIF on average over 10 runs. Column REUSED reports the MS obtained after reusing inputs to kill additional mutants. Column DIFF shows the differences between FUZZED and REUSED, and reports on the significance of the difference (we highlight significant improvements). The next three sets of columns report statistics (mean, min, and max) on the number of inputs available for reuse across target functions, the time required to test each mutant with the reused inputs, and the time taken to kill a mutant. Regarding the execution time, we report statistics based on the 500 execution results obtained as described in Section 5.7.1.
The experiment results show that, in the best case (i.e., LIBU), the MS improves by 3.79 pp and, in the worst case (i.e., ESAIL), it improves by 0.12 pp. The average improvement across the five subjects is 1.40 pp. Since in most of the subjects the improvement is above 1 pp, we conclude that input reuse is beneficial for MOTIF. The time taken for reusing inputs varies, taking up to 15 minutes (900 seconds) depending on several factors, including the number of inputs, software size, and the type of mutants. For instance, ESAIL has approximately 70% of its mutants within a single function, which leads to a large number of additional executions when reusing inputs. In the case of S5, the larger size causes a longer execution time for fuzzing drivers compared to other subjects, but the maximum execution time is lower due to the number of inputs collected. The execution time is also affected by the type of mutants. Specifically, some mutants took longer than others because they modified stopping conditions in program loops.
Table 10: Comparison MOTIF with Best set of fuzzing options and results after reusing inputs for five selected subjects.
Subject Target mutants 3 c MS (Avg. of 10 runs) 3 c # of inputs for functions (10 runs) 3 c Reuse time (10$\times$50 runs, seconds) 3 c Kill time (10$\times$50 runs, seconds)
@@ 5.7.2 Results
n.nn pp: outperformed the baseline with $p\leq 0.01$
@@
6 Threats to validity
We discuss internal, conclusion, construct, and external validity according to standard practice wohlin2024experimentation ; feldt2010validity ; Ralph2018 .
6.1 Internal validity.
@@ 6 Threats to validity
Additionally, a potential threat may arise due to our results depending on the specific fuzzer used in our experiments, AFL++. However, AFL++ is one of the best performing grey-box fuzzers according to recent benchmarks (see Section 5.2). Further, we assessed MOTIF with a state-of-the art hybrid fuzzing solution. Although alternative hybrid fuzzers were not applicable to some of our subjects (see Section 5.2), they might have led to different and potentially better results; we leave their investigation to future work.
6.2 Conclusion validity.
Threats to conclusion validity relate to statistical power of our results, invalid statistical test assumptions, reliability of measurements, and random irrelevancies  wohlin2024experimentation .
To deal with non-determinism of the different tools considered in our experiments (MOTIF, SEMuP, MOTIF-Hybrid), we repeated the execution of each tool ten times. To discuss the significance of the differences across tools, we applied the non-parametric Mann Whitney U-test because the distribution of mutants killed is not known; U-test is also a common choice in software engineering papers Arcuri:practicalGuide:ICSE:2015 .
Our measurements concern the number of killed mutants and execution time; for execution time we relied on the timing functions provided by the operating system, while for mutants killing we verified the correctness of our tools. We do not expect unreliable measurements.
In our context, the only source of random noise might be the workload of the machines used to run the experiments, which may slow down the performance of some of the tools. To mitigate this threat, in addition to relying on high-performance computing nodes with guarantees for the provided service level, we manually inspected the execution logs to ensure that no anomalies — such as exceptions caused by the host environment — were present.
6.3 Construct validity.
Construct validity concerns face, content, convergent, and predictive validity Ralph2018 . The constructs considered in our work are effectiveness and cost. Effectiveness is measured through the percentage of live mutants killed, while cost is measured in terms of execution time.
Face validity concerns the selection of appropriate reflective indicators. For effectiveness, we rely on the percentage of live mutants killed, which is appropriate in this context because the final objective of the tools considered in our experiments is to maximize the percentage of killed mutants. For cost, we measured the time required to kill mutants; specifically, for RQ1 to RQ4 we set a fixed time budged for all the tools, for RQ5 we measured the time taken to re-execute all the inputs and the time taken to kill mutants. The time required to kill mutants is an appropriate measure of cost because it indicates how long engineers have to wait before improving their test suite and has an impact on the whole testing process.
Content validity concerns the adequacy of reflective indicators to cover the breadth of the construct. For effectiveness, we could consider the fault detection rate, measured as the proportion of real-world faults detected by the test cases generated by the different tools. However, fault detection rate can be significantly assessed only with a large set of real-world faults, which we do not have access to, for our subjects. For cost, we could consider the time required to configure the assessed tools, which we did not directly measure. To mitigate this threat, we report that RQ1 is not affected by such issue because all the tools have the same configuration time since they differ only regarding the selection of configuration parameters. In RQ2, we minimized the threat by constructing a pipeline for symbolic execution that resembles MOTIF’s. RQ3 and RQ5 do concern the comparison of tools with different configurations. For RQ4, we report that hybrid fuzzing has a slightly more complex setup.
We are not affected by convergence validity threats because we collected only one reflective indicator per construct. To address predictive validity, we reported on significance of the differences when comparing different tools (RQ1, RQ2, RQ4).
6.4 External validity.
@@
Threats to external validity may arise from the use of a limited set of subjects for our experiments, which may affect the generalizability of our results. To address this issue, we selected diverse software subjects that are installed and running on space CPS, including satellites currently in orbit: a mathematical library, a utility library, a data serialization component, onboard control software, and ground software. Since they implement a diverse set of features (mathematical operations, serialization, string and time utilities), they strengthen the generalizability of our results. Further, these types of software components are typical in many CPS systems including avionics, robotics, and automotive, thus suggesting the proposed approach may be useful in many sectors other than space.
Another threat to external validity comes from the nature of the mutants considered in our experiments, which are constructed by MASS using sufficient and deletion operators. Note that these operators are the ones commonly implemented in other state-of-the-art mutation analysis tools papadakis2019mutation . Further, they were considered in experiments demonstrating correlation between mutation score and fault detection capabilities of test suites papadakis2018mutation ; Chekam:17 ; consequently, test cases killing those mutants likely detect other mutants. We leave to future work the investigation of MOTIF effectiveness for other types of mutants, such as mutants generated using generative AI muBert , mutants focusing on cybersecurity properties of software muVuln , or mutants altering characteristics specific for object-oriented programs delgado2017assessment .